The sum of the digits of a two-digit number is 11.Twice the tens digit plus 2 equals ten times the one digit. What is the number?
Answer by swincher4391(1107)

x+y = 11
2x + 2 = 10y
y = 11-x
2x + 2 = 10(11-x)
x+1 = 5(11-x)
x+1 = 55 - 5x
6x = 54
x = 9
y = 2
The number is 92.